A chemist has to mix two chemicals, A and B, to make 150 grams of a solution. This solution requires 10 grams of A or every gram of B. If the amount of chemical A is designated as x and the amount of Chemical B is designated y, what equation would you use to find out how much of each chemical she will need.

Answer by jorel1380(3719)

a=10b
a+b=150
x+y=150
x=10y
10y+y=150
y=150/11
x=1500/11..